Factor 80 into a product of prime factors.
Answer by stanbon(48569)

Factor 80 into a product of prime factors
Procedure.
Divide by the smallest possible prime as many times as you can:
80/2 = 40
40/2 = 20
20/2 = 10
10/2 = 5
---------
Then 80 = 2^4*5
==================
Cheers,
Stan H.